OpenStudy (anonymous):

Determine whether the proportion is true or false. 35/25= 7/5

OpenStudy (unklerhaukus):

true as

OpenStudy (anonymous):

(7*5)/(5*5) =35/25 or 7/5 = (35/5)/(25/5) thus 7/5=7/5 yes

OpenStudy (anonymous):

simplify by 5: 35/25 = 7/5

OpenStudy (anonymous):

Final Answer is TRUE